# Schedule Your Edit: A Simple yet Effective Diffusion Noise Schedule for Image Editing

### v1

@@ 1 Introduction
Text-guided diffusion models have emerged as a leading technique in image generation, offering remarkable visual quality and diversity [2, 42, 50, 69]. The noise latent space of these models can be leveraged to retain and modify images [32, 66, 68], enabling text-guided editing where a source image is adjusted based on a target prompt. This requires first inverting the source image into a latent variable (e.g., via DDIM inversion), due to the absence of its predefined latent space [28, 39].
@@ 4.1 Well-Defined Schedule Improve Inversion Stability
where $k$ and $t$ are hyperparameters that control the steepness and midpoint of the logistic function, respectively. In our experiments, we set $k=0.015$ and $t_{0}=\text{int}(0.3T)$, as discussed in Section 5.3.1.
@@ 4.2 Exploring Noise Space of Logistic Schedule
Comprehensive Pattern Capture. As shown in Fig. 4 (right), our logistic schedule preserves more original image information during the initial stages of inversion, reflecting the logSNR trend. Moreover, the steady denoising process ensures that the inverted latent, especially in the final stages, closely resembles pure Gaussian noise, minimizing the retention of low-frequency components. This thorough process ensures that the inversion encodes a broader array of the original image’s information, thereby enhancing the quality and fidelity of the edited images.
@@ 5.3 Ablation Studies
In this section, we investigate the effects of different configurations of the Logistic Schedule and the adaptability of the Logistic Schedule with various inversion techniques and diffusion models. More experiments on hyperparameters (e.g., guidance scale, input scale) can be found in Appendix E.
@@ 5.3.1 Effects of Configuration of Logistic Schedule
We conduct experiments with different configurations of the Logistic Schedule in Eq. 5, providing further evidence for the noise space analysis (Section 4.2).
Different $k$: Changing the shape of logSNR. When $k$ is larger, the logSNR values span a larger range (Fig.6, left). However, if the range is too large, it results in excessive loss of original image information in edited images (Fig.6, right). Interestingly, when $k$ is small, the logSNR resembles that of linear and cosine schedules, but the logistic schedule better preserves the original image content without altering the overall structure. This further supports Proposition 3.1 that the singularity in linear and cosine schedules tends to destroy original image information, causing undesired changes.
@@ 5.3.2 Adapting Inversion Techniques and Diffusion Models
To validate the adaptability and robustness of the Logistic Schedule, we first apply it with Plug-and-Play [59] using Stable Diffusion v1.5 [50] as the baseline. We then design experiments with two other diffusion models, Stable Diffusion v2.1 and Stable Diffusion XL [47], and incorporate three advanced inversion approaches: Null-Text Inversion [41], Negative Prompt Inversion [40], and Direct Inversion [25]. As shown in Table 2, while more advanced stable diffusion models increase textual similarity, they degrade content preservation. Conversely, incorporating advanced inversion approaches improves both content preservation and edit fidelity.
@@ 6 Conclusion
This paper introduces Logistic Schedule, a novel noise schedule designed to eliminate singularities and improve inversion stability for image editing, thereby opening new avenues for noise schedule designs in image editing tasks. By reducing noise prediction errors, Logistic Schedule enhances content preservation and edit fidelity. Extensive experiments demonstrate its superior performance and adaptability across various image editing tasks compared to traditional noise schedules.
@@ B.1.2 Cosine Schedule
B.1.3 Logistic Schedule
@@
$\bar{\alpha}_{t}=\dfrac{1}{1+e^{-k(t-t_{0})}},$
@@ Appendix E Experimental Results
E.1 Effects of Guidance Scale
E.2 Quantitative Comparison Across Editing Types
@@
E.3 Qualitative Comparison Across Editing Types
The logistic noise schedule consistently outperforms linear and cosine schedules across various editing tasks. It excels in preserving the original image content while making specified changes without artifacts (Fig. 11), maintaining natural and coherent lighting in color edits (Fig. 12), and accurately rendering new material properties (Fig. 13). For object switching and addition, it ensures seamless integration with consistent lighting and spatial relationships (Figs. 14, 15). In non-rigid editing, it preserves anatomical correctness and smooth transitions (Fig. 16). It also blends new backgrounds naturally in scene transfers (Fig. 17) and maintains essential details while applying new artistic styles (Fig. 18).
E.4 Effects of Input Scale
Input Scale Structure Background Preservation CLIP Similarity (%)
w/o Normalizing $b$ 22.40 22.91 90.75 82.05 79.32 86.72 22.65
w. Normalizing $b$ 24.46 22.06 108.43 83.47 79.16 86.24 22.48
@@ Appendix H Ethics Statement
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The main claims made in the abstract and introduction (Section 1) accurately reflect the paper’s contributions and scope.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The limitations of the work are discussed in Appendix F.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Concerning Proposition 3.1, Appendix B provides its corresponding proofs, with a short proof sketch to provide intuition.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Refer to Appendix D.1 and D.2 for implementation details, the provided code in the supplementary materials ensures the reproducibility as well.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Refer to the provided code in the supplementary materials.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Table 3 specifies all the tasks included in this work. Appendix D.2 provides the employed hyperparameters.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Appendix E.2 provides the statistical significance of the experiments.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Appendix D.2 provide sufficient information on the computer resources.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The authors have reviewed and understand the NeurIPS Code of Ethics, and confirm that their research conforms to it in every respect.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Potential societal impacts of the work are discussed in Appendix G.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The paper poses no such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The original papers of assets, and the corresponding versions are provided in Section 5, References and Appendix D.2.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We will document the assets well when we officially release our code.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not describe potential risks incurred by study participants.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.

### v3

@@ 1 Introduction
Text-guided diffusion models have emerged as a leading technique in image generation, offering remarkable visual quality and diversity [2, 42, 50, 69] This work was completed during the internship at SGIT AI Lab, State Grid Corporation of China. . The noise latent space of these models can be leveraged to retain and modify images [32, 66, 68], enabling text-guided editing where a source image is adjusted based on a target prompt. This requires first inverting the source image into a latent variable (e.g., via DDIM inversion), due to the absence of its predefined latent space [28, 39].
@@ 4.1 Well-Defined Schedule Improve Inversion Stability
where $k$ and $t$ are hyperparameters that control the steepness and midpoint of the logistic function, respectively. In our experiments, we set $k=0.015$ and $t_{0}=\text{int}(0.6T)$, as discussed in Section 5.3.1.
@@ 4.2 Exploring Noise Space of Logistic Schedule
Comprehensive Pattern Capture. As shown in Fig. 4 (right), we visualize the latents during the inversion (forward) process, using 50 timesteps with the final step at 981 instead of 999. In the early stage, our Logistic Schedule preserves more original image information, reflecting the logSNR trend. Considering the later stage, linear and cosine schedules retain more low-frequency components due to higher endpoint SNRs, explaining why their noise maps don’t fully cover the image. In contrast, our Logistic Schedule ensures that the inverted latent closely resembles pure Gaussian noise, minimizing the retention of low-frequency components. This thorough process ensures that the inversion encodes a broader array of the original image’s information, thereby enhancing the quality and fidelity of the edited images.
@@ 5.3 Ablation Studies
In this section, we investigate the effects of different configurations of the Logistic Schedule and the adaptability of the Logistic Schedule with various inversion techniques and diffusion models. More experiments on hyperparameters (e.g., guidance scale, input scale) can be found in Appendix E. The comparison with more design of the noise scheduler is provided in Appendix E.4.
@@ 5.3.1 Effects of Configuration of Logistic Schedule
We conduct experiments with different configurations of the Logistic Schedule in Eq. 5, providing further evidence for the noise space analysis (Section 4.2). The parameters of the Logistic Schedule (Eq. 5)—specifically the steepness ($k$) and the midpoint ($t_{0}$)—play a crucial role in balancing content preservation and edit fidelity. Table 2 provides the quantitative results of varying $k$ and $t_{0}$.
Table 2: Quantitative results of the Logistic Schedule across various hyperparameter settings. The best method is indicated in bold, and the worst method is shown in purple.
Settings Structure Background Preservation CLIP Similarity (%)
$k=0.015$ $t_{0}=\text{int}(0.6T)$ 17.37 24.78 81.80 49.47 82.97 82.44 23.62
$k=0.008$ 16.27 26.45 75.62 43.20 85.28 84.07 20.47
$k=0.011$ 16.64 25.80 77.90 48.83 84.15 83.76 21.46
$k=0.017$ 22.79 22.33 99.98 57.32 81.52 82.10 23.25
$k=0.029$ 27.82 21.05 103.45 64.36 78.48 80.66 23.81
$t_{0}=\text{int}(0.4T)$ 24.31 22.41 97.21 60.84 79.72 79.47 20.33
$t_{0}=\text{int}(0.8T)$ 29.47 21.64 95.58 63.89 75.14 77.05 22.68
Different $k$: Changing the steepness of logSNR. When $k$ is larger, the logSNR values span a larger range (Fig.6, left). However, if the range is too large, excessive steepness of logSNR results in excessive loss of original image information in edited images (Fig.6, right). Interestingly, when $k$ is small, the logSNR resembles that of linear and cosine schedules, but the logistic schedule better preserves the original image content without altering the overall structure. This further supports Proposition 3.1 that the singularity in linear and cosine schedules tends to destroy original image information, causing undesired changes.
@@
Balancing these parameters, we find that $k=0.015$ and $t_{0}=\text{int}(0.6T)$ strike the optimal trade-off between content preservation and edit fidelity, providing robust performance across various tasks.
@@ 5.3.2 Adapting Inversion Techniques and Diffusion Models
To validate the adaptability and robustness of the Logistic Schedule, we first apply it with Plug-and-Play [59] using Stable Diffusion v1.5 [50] as the baseline. We then design experiments with two other diffusion models, Stable Diffusion v2.1 and Stable Diffusion XL [47], and incorporate three advanced inversion approaches: Null-Text Inversion [41], Negative Prompt Inversion (NPI) [40], and Direct Inversion [25]. As shown in Table 3, while more advanced stable diffusion models increase textual similarity, they degrade content preservation. Conversely, incorporating advanced inversion approaches improves both content preservation and edit fidelity.
@@
Furthermore, Table 4 presents the detailed comparison between the Logistic Schedule and the scaled linear schedule across different inversion techniques.
Table 4: Comparison of inversion techniques with the scaled linear schedule and our proposed Logistic Schedule.
Variants Structure Background Preservation CLIP Similarity
Dist $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ MSE $\downarrow$ SSIM $\uparrow$ Visual $\uparrow$ Textual $\uparrow$
Null-Text + Linear 21.00 23.00 95.00 63.00 81.50 81.00 21.30
Null-Text + Logistic 18.67 23.80 89.64 57.97 82.97 82.46 21.95
NPI + Linear 28.00 22.40 105.00 78.00 78.50 78.50 20.90
NPI + Logistic 24.82 23.17 99.19 71.24 80.26 80.16 21.53
Direct + Linear 19.00 24.90 78.00 45.00 83.50 82.90 21.90
Direct + Logistic 16.06 25.73 74.17 41.19 85.61 83.29 22.03
@@ 6 Conclusion
This paper presents the Logistic Schedule, a novel noise schedule that eliminates singularities and improves inversion stability for image editing. Our method enhances content preservation and edit fidelity without requiring additional retraining, making it a plug-and-play solution for existing workflows. Through in-depth analysis of the diffusion inversion process, we identify that current schedulers suffer from singularity issues at the start of inversion. The proposed Logistic Schedule provides a straightforward solution to this problem, offering superior performance and adaptability across various image editing tasks.
7 Acknowledgement
This work was supported by National Natural Science Foundation of China (62293551, 62377038,62177038,62277042). Project of China Knowledge Centre for Engineering Science and Technology, Project of Chinese academy of engineering “The Online and Offline Mixed Educational Service System for ‘The Belt and Road’ Training in MOOC China". “LENOVO-XJTU" Intelligent Industry Joint Laboratory Project.
@@ B.1.2 Cosine Schedule
B.1.3 Sigmoid Schedule
The sigmoid schedule is introduced in Jabri et al., which is designed for scalable data generation, especially for high-dimensional data, without addressing the challenges of DDIM inversion. The formulation of the sigmoid schedule can be presented as below:
$\tilde{\alpha}_{t}=\dfrac{-\left(\frac{t(e-s)+s}{r}\right)\cdot\text{sigmoid}()+v_{e}}{v_{e}-v_{s}}$ (21)
where $s$ and $e$ are the start and end of the sigmoid function’s range, and $v_{s}=(s/r)\cdot\text{sigmoid}(),\;v_{e}=(e/r)\cdot\text{sigmoid}()$.
B.1.4 Logistic Schedule
@@
$\bar{\alpha}_{t}=\text{Normalized}\left(\dfrac{1}{1+e^{-k(t-t_{0})}}\right),$
@@ B.2.2 Cosine Schedule
B.2.3 Sigmoid Schedule
Proof.
Given definition of $x_{t}$:
$\mathbf{x}_{t}=\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon_{t}$
To express the coefficient of $\epsilon$ in the derivative of $\mathbf{x}_{t}$ with respect to $t$, we start with the expression:
$\mathbf{x}_{t}=a(t)\cdot\epsilon+b(t)\cdot\mathbf{x}_{0},$
where $\epsilon$ represents the noise, and $\mathbf{x}_{0}$ is the original image. Given that $\epsilon$ and $\mathbf{x}_{0}$ are constants with respect to $t$, the differentiation yields:
$\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{x}_{t}=\epsilon\cdot\frac{\mathrm{d}}{\mathrm{d}t}a(t)+\mathbf{x}_{0}\cdot\frac{\mathrm{d}}{\mathrm{d}t}b(t).$ (24)
Recall the definition of $\bar{\alpha}_{t}$ in sigmoid schedule in Eq. 21, we put it in the expression of $a(t)$, then the coefficient of $\epsilon$ in Eq. 24 can be expressed as:
$\frac{\mathrm{d}}{\mathrm{d}t}a(t)=\frac{(e-s)e^{-s-t(e-s)}}{2\tau}\left(\sqrt{1-\left(\frac{\frac{1}{1+e^{\frac{s+t(e-s)}{\tau}}}-\frac{1}{1+e^{\frac{t(e-s)}{\tau}}}}{\frac{1}{1+e^{\frac{s}{\tau}}}-\frac{1}{1+e^{\frac{t(e-s)}{\tau}}}}\right)^{2}\left(e^{-s-t(e-s)}+1\right)^{2}}\right)^{-1}$
When $t\rightarrow 0$, we have the derivative diverges to infinity:
$\lim_{t\to 0}\frac{\mathrm{d}}{\mathrm{d}t}a(t)\rightarrow\infty,$
@@ B.3 Derivatives of the Logistic Schedule
B.4 Clarification of Differences and Motivations
Motivations of the Logistic Schedule: The primary focus of the Logistic Schedule is on addressing the singularity issues in DDIM forward (inversion) processes, enhancing both inversion stability and noise space.
@@ Appendix E Experimental Results
E.1 Quantitative Comparison Across Editing Types
@@
E.2 Qualitative Comparison Across Editing Types
The logistic noise schedule consistently outperforms linear and cosine schedules across various editing tasks. It excels in preserving the original image content while making specified changes without artifacts (Fig. 13), maintaining natural and coherent lighting in color edits (Fig. 14), and accurately rendering new material properties (Fig. 15). For object switching and addition, it ensures seamless integration with consistent lighting and spatial relationships (Figs. 16, 17). In non-rigid editing, it preserves anatomical correctness and smooth transitions (Fig. 18). It also blends new backgrounds naturally in scene transfers (Fig. 19) and maintains essential details while applying new artistic styles (Fig. 20).
E.3 Broader Application: Training-Based Methods
While the Logistic Schedule has shown broad applicability in various image editing tasks, we further explore its use in training-based methods, specifically in Text-to-Image synthesis. For this, we conducted experiments by fine-tuning the UNet using DreamBooth [51], leveraging approximately 100 images for each configuration.
Table 7: Performance comparison of DreamBooth fine-tuning using different noise schedules on the SD-1.5 model. The best results are in bold.
Setting DINO ($\uparrow$) CLIP-I ($\uparrow$) CLIP-T ($\uparrow$) PSNR ($\uparrow$) PRES ($\uparrow$)
Real Images 0.823 0.902 N/A 26.86 0.653
DreamBooth (SD-1.5) + Linear 0.726 0.823 0.268 24.41 0.567
DreamBooth (SD-1.5) + Cosine 0.745 0.857 0.264 24.75 0.554
DreamBooth (SD-1.5) + Logistic 0.761 0.877 0.293 25.62 0.580
The related qualitative comparisons are shown in Fig. 9. These results demonstrate that training DreamBooth with the Logistic Schedule improves performance across key metrics such as DINO and CLIP-I similarity, as well as PSNR and PRES, outperforming both linear and cosine schedules.
Figure 9: Qualitative comparisons of fine-tuning DreamBooth using different noise schedules (Linear, Cosine, and Logistic). The top column presents the fine-tune samples, and the instruction prompts, and the below column displays the corresponding fine-tuned outputs. The Logistic Schedule produces superior outputs with improved fidelity and alignment to the prompts.
E.4 Comparison With Other Noise Schedulers
We conducted experiments comparing our Logistic Schedule with other schedules under the DDIM paradigm, such as exponential, sigmoid, hyperbolic, and geometric schedules. Table 8 displays the quantitative results. The best-performing method is indicated in bold, the worst method is marked in purple, and the second-best method is underlined.
Table 8: Comparison of different noise schedules. Metrics include Structure Distance (×$\times 10^{-3}$), PSNR (higher is better), LPIPS (×$\times 10^{-3}$, lower is better), MSE (×$\times 10^{-4}$, lower is better), SSIM (×$\times 10^{-2}$, higher is better), and CLIP Similarity for both visual and textual content.
Schedule Dist $\downarrow$ PSNR $\uparrow$ LPIPS $\downarrow$ MSE $\downarrow$ SSIM $\uparrow$ Visual $\uparrow$ Textual $\uparrow$
Linear 35.66 20.70 134.88 113.61 77.60 79.82 23.06
Cosine 26.57 22.38 110.52 80.01 80.15 81.35 22.39
Exponential 16.22 25.20 80.45 47.11 82.23 82.78 19.50
Hyperbolic 36.55 20.95 140.55 119.78 79.89 79.20 23.20
Geometric 18.12 24.10 92.45 62.13 82.05 82.20 20.45
Sigmoid 27.80 22.55 115.32 85.60 80.22 81.50 22.55
Logistic 17.37 24.78 81.80 49.47 82.97 82.44 23.62
Figure 10: Qualitative comparison between different noise schedules for both reconstruction and editing tasks.
The results demonstrate that our Logistic Schedule achieves competitive performance across various metrics. Notably, it offers significant improvements in content preservation and edit fidelity compared to other schedules. As shown in Fig. 10, the Logistic Schedule preserves the visual characteristics of the source image more faithfully during reconstruction and enables more precise control during editing.
@@ E.5 Reconstruction Ability of Different Noise Schedule
E.6 Effects of Guidance Scale
E.7 Effects of Input Scale
Input Scale Structure Background Preservation CLIP Similarity (%)
w/o Normalizing $b$ 22.40 22.91 90.75 82.05 79.32 86.72 22.65
w. Normalizing $b$ 24.46 22.06 108.43 83.47 79.16 86.24 22.48